in a hexagon three angles have the same measure. The measure of each of the congruent angles is twice the measure of the fourth angle and is half the measure of the fifth angle. The sixth angle measures 115. What is the measure of the smallest angle?
Answer by gonzo(654)

a hexagon has 6 angles.
the total number of internal degrees is (n-2)*180 = 4*180 = 720 degrees.
-----
let a = one of the congruent angles.
you have:
angles 1, 2, and 3 = a.
-----
if a = 2 * angle 4, then angle 4 = 1/2 * a
-----
if a = 1/2 * angle 5, then angle 5 = 2 * a
-----
we have:
angle 1 = a
angle 2 = a
angle 3 = a
angle 4 = 1/2 * a
angle 5 = 2 * a
angle 6 = 115
-----
since the sum of the angles is equal to 720, we have:
3*a + 2*a + 1/2*a + 115 = 720
combine like terms to get:
5.5 * a + 115 = 720
subtract 115 from both sides to get:
5.5 * a = 720 - 115 = 605
divide both sides by 5.5 to get:
a = 720/5.5 = 110
-----
angle 1, 2, and 3 are each 110 degrees.
angle 4 = 55 degrees
angle 5 = 220 degrees
angle 6 = 115 degrees
-----
the question is:
What is the measure of the smallest angle?
the answer is:
55 degrees.